Return on Common Stock You buy a share of The Ludwig Corporation stock for $23.10. You expect it to pay dividends of $1.06, $1.1427, and $1.2318 in Years 1, 2, and 3, respectively, and you expect to sell it at a price of $28.94 at the end of 3 years.

A). Calculate the growth rate in dividends. Round your answer to two decimal places.

B). Calculate the expected dividend yield. Round your answer to two decimal places.

C). Assuming that the calculated growth rate is expected to continue, you can add the dividend yield to the expected growth rate to obtain the expected total rate of return. What is this stock's expected total rate of return (assume market is in equilibrium with the required rate of return equal to the expected return)? Do not round intermediate calculations. Round your answer to two decimal places. %
Answer #1

Growth rate in dividends = (Dividend in Year 2 - Dividend in Year 1)/Dividend in Year 1

= (1.1427-1.06)/1.06

= 7.80%

In year 3 = (1.2318 - 1.1427)/1.1427= 7.80%

Expected Dividend yield = Dividend in Year 1/Current Stock price

= 1.06/23.10

= 4.59%

Total rate of return = 7.80%+4.59%

= 12.39%